A 5kg block was placed on an inclined plane(50 degrees with respect to the vertical).Forces A is at angle C with respect to the horizontal. Force B is parallel with the incline.The coefficients of static and kenetic friction are 0.3 and 0.2 respectively.

A=197KN

B=49 KN

C=7 degrees

A.DETERMINE IF THERE IS MOTION AND ITS DIRECTION

B.BASED ON YOUR ANSWER IN "A" WHAT IS THE ACCELERATION OF THE BLOCK?


1
Expert's answer
2020-12-31T08:07:02-0500

A.


"N=mg\cos40°+197000\cdot\sin47°=5\cdot9.8\cdot\cos40°+197000\cdot\sin47°="


"=144114(N)"


"49000+mg\sin40°-197000\cos47°=49000+5\cdot9.8\cdot\sin40°-197000\cos47°="


"=-85322(N)"


"F_f=0.3N=0.3\cdot144114=43234(N)"


"F_{total}=-85322+43234=-42088(N)"


The block will move up the inclined plane. Answer


B.


"a=F\/m=42088\/5=8418(m\/s^2)" . Answer